# Aerial Vision-and-Language Navigation via Semantic-Topo-Metric Representation Guided LLM Reasoning

### v1

@@ III-C Designing Text Prompt for LLM Planning
1. (Completed) Lift off 5 meters.
2. (Completed) Turn 180 degrees.
3. (Completed) Head straight across the water to the road.
4. (In Process) Turn left to align the road.
@@
IV Experiment Setup
TABLE I: Comparison on the validation seen and unseen set of the AerialVLN-S dataset.
Validation Seen Validation Unseen
# AerialVLN-S NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$
S1 Random 109.6 0.0 0.0 149.7 0.0 0.0
S2 Action Sampling 213.8 0.9 5.7 237.6 0.2 1.1
S3 LingUNet [misra2018mapping] 383.8 0.6 6.9 368.4 0.4 3.6
S4 Seq2Seq [anderson2018vision] 146.0 4.8 19.8 218.9 2.3 11.7
S5 CMA [liu_2023_AerialVLN] 121.0 3.0 23.2 172.1 3.2 16.0
S6 LAG [liu_2023_AerialVLN] 90.2 7.2 15.7 127.9 5.1 10.5
S7 Ours 96.3 12.6 31.6 119.5 10.8 23.0
Datasets
We conduct experiments using a novel and complex dataset provided by [liu_2023_AerialVLN], which we refer to as the "AerialVLN-S" dataset, to assess the performance of aerial VLN tasks. This dataset is designed to mimic real-world urban environments with more than 870 different kinds of objects and various scenarios including downtown cities, factories, parks, and villages. Besides, the AerialVLN-S dataset encompasses 8,446 flying paths recorded by experienced human UAV pilots who possess the AOPA (Aircraft Owners and Pilots Association) certificate, ensuring the paths’ realism and quality. Several comparison methods are evaluated on this dataset, including state-of-the-art VLN models, to establish a benchmark for aerial VLN tasks. We utilize AerialVLN-S as the benchmark, with 12 scenes from the validation seen set and 5 scenes from the validation unseen set, respectively. Notably, the test set is not used because it is inaccessible and the evaluation server [liu_2023_AerialVLN] is unavailable now.
Evaluation metrics
@@ IV Experiment Setup
To rigorously assess the performance of our method, we utilize a comprehensive set of evaluation metrics as in [liu_2023_AerialVLN]. We focus on several key aspects, i.e., Navigation Error (NE) quantifying the distance between the UAV’s stopping point and the actual destination, Success Rate (SR) which measures the proportion of navigations that successfully reach the destination within a 20-meter threshold, and Oracle Success Rate (OSR), an idealized measure considering any point on the trajectory that comes within 20 meters of the destination as a success.
Implementation details
@@
Our framework is implemented in both a simulator and a real environment. For the simulation environment, we evaluated the algorithm on airsim and UE4, on a laptop with an Intel i9 12th generation CPU and an Nvidia RTX 4090 GPU. For the outdoor environment, we test on a Q250 airframe, carrying an Intel RealSense D435i depth camera, and an NVIDIA Jetson Xavier NX running Ubuntu 18.04 as onboard computer. We use the online API of GPT-4V and GPT-4o, where the default parameters are employed.
IV-A Experimental Results
Quantitative results in simulator
@@ IV-A Experimental Results
Table I compares our method with several other works. Random and Action Sampling respectively mean sampling actions randomly and according to the action distribution of the training set. LingUNet [misra2018mapping] assumes an agent can see the destination from the beginning, which is transformed into a step-wise paradigm by [liu_2023_AerialVLN]. Seq2Seq [anderson2018vision] is a learning-based recurrent policy, directly predicting the next action. CMA is a baseline model in [liu_2023_AerialVLN], considering attention between multi-modalities. LAG is the main method in [liu_2023_AerialVLN], designing a look-ahead guidance to adjust paths. In comparison, our method significantly outperforms others in most metrics, demonstrating the superiority of our method. We found that instead of processing the entire long instruction and visual observations at each time step, the formatted decomposed sub-goal is more conducive to the fine-grained alignment of each action and landmark. Moreover, UAV has a better understanding of its own spatial position and other semantic constraints under the guidance of our STMR, which has obvious advantages in long instruction environments and leads to more robust execution.
Quantitative results in real environment
@@
To test whether our solution works well in the real world, we collect 10 outdoor scenes including street scenes and forests with ground-truth lengths ranging from 50m to 500m. Then, we apply our method on real UAVs and call cloud-hosted LLMs (GPT-4o is used) to navigate in these challenging environments. LLM-based VLN methods MapGPT [chen2024mapgpt] and NavGPT [zhou2024navgpt] are employed and compared in this experiment. From demo snapshot Fig. 4, we can see that our method can align visual and textual landmarks as well as understand commands. Finally, the UAV reaches the destination successfully.  Table II also shows that the proposed method achieves much better results than most recent LLM-based VLN methods.
@@
Figure 4: Qualitative results in real environment. Short-range movements of the same action are merged into a single long-range movement, as shown in the images sequence.
TABLE II: Qualitative results in real environment.
Module SR/$\uparrow$ OSR/$\uparrow$
MapGPT [chen2024mapgpt] 20 20
NavGPT [zhou2024navgpt] 10 20
Ours 40 70
Failure case analysis
@@
In the simulation, we analyze two most common failures caused by incorrect planning or execution. The first typical failure is the misunderstanding of ambiguous instructions. In some instructions, there are a lot of continuous commands without any landmark, such as "turn left, then move right, then go straight", which have no objects for an LLM to refer to. This case will directly leads to the repeated execution of a step. The second typical error is caused by the inaccuracy of the visual perceptor. Although modern perception models show strong ability, they still need to be improved in identifying objects from different views. As a result, the key landmark may not be mapped to the matrix representation, causing the proper action not to be completed. This observation aligns with our real-world experiment, where most errors are from perceptions.
TABLE III: Ablation study on randomly selected 100 scenes from Validation Unseen dataset for different semantic spatial representations as prompts.
# Method Validation Unseen
NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$
1 Topo 203.3 4.9 12.8
2 Metric 165.0 6.1 13.7
3 Ours 87.1 19.4 26.2
IV-B Comprehensive Analysis
@@
We carry out several ablation studies to assess the effectiveness of the proposed framework. All ablation experiments are conducted on the AerialVLN-S task with the validation unseen dataset. Corresponding results are depicted in Table III,  Table IV and Table V.
Explicit spatial information
@@
To demonstrate the spatial representation ability of STMR, we further compare it with other prompting formats as an LLM’s spatial information prompt. In Topo format (row 1 of Table III, we maintain a linguistic-formed map that captures the topological relationships between different nodes. Each node records textual descriptions of visual observations, and the connectivity between nodes is described using textual prompts. For example, "Place 1 is connected with Places 2, 4, 0, 3". In Metric format (row 2 of Table III, we arrange the visual observations from 8 different directions in a clockwise order relative to the agent’s current orientation and concatenate them into a single prompt. Each vision observation records the direction and distance of the landmark, for example, "a white building in the left front 10 meters away". It is shown that row 2 improves SR by 1.2% (absolute) over row 1 but still poorer navigation performance, since for some landmarks such as roads and rivers, UAVs still cannot obtain region-level spatial information based on a certain point distance description. Our STMR combining the semantic, topological, and metric information significantly enhances the UAV’s exploration capability and improves the accuracy in SR and OSR by over 10%.
State update of sub-goal
To ensure that each sub-goal will be executed despite environment changes, in each iteration, the LLM does not change the original path plan, but updates the status of the current sub-goal according to the observation, including todo, in process, and completed. We compare this strategy with updating the entire multi-step path planning (row 1) according to the observation in each iteration. The results are shown in Table IV. This strategy does not notably enhance the Oracle Success Rate (OSR), showing that the proposed STMR already furnishes enough capacity for navigating the environment comprehensively. Nonetheless, it positively impacts the decision-making process, resulting in an increase in the success rate (SR) from 9.0% to 10.8%.
TABLE IV: Ablation study for the state update.
Module NE/$\downarrow$ SR/$\uparrow$ OSR/$\uparrow$
w/o state 289.7 9.0 22.1
w state 219.8 10.8 23.0
TABLE V: Ablation study for the direct use of RGB images and the proposed STMR as the visual prompt for LLMs.
LLM Input RGB STMR NE/$\downarrow$ SR/$\uparrow$ OSR/$\uparrow$
GPT-4V $\checkmark$ 350.0 1.5 9.7
$\checkmark$ 193.1 5.6 17.4
GPT-4o $\checkmark$ 412.0 1.1 10.0
$\checkmark$ 119.5 10.8 23.0
Direct visual input and STMR
@@
Due to the strong multi-modal capabilities of GPT-4V and GPT-4o, they can reason with both visual and textual prompts. However, it is challenging for them to understand landmark semantics directly from step-by-step input RGB images. This may be because the inability of LLMs in spatial-aware reasoning. As shown in Table V, the proposed STMR representation with LLMs significantly outperforms the direct input of RGB images with VLMs.
@@ V CONCLUSIONS AND FUTURE WORK
This paper addresses the challenging aerial VLN task by introducing an LLM-based end-to-end framework. To enhance the spatial reasoning ability of LLMs, we propose Semantic-Topo-Metric Representation (STMR). STMR first integrates instruction-related landmarks and their locations into a top-down map, and subsequently transforms this map into a matrix representation containing semantic, topological, and distance metric information. Taking the proposed STMR as a part of prompts of the LLM, we significantly improve the UAV’s navigation capabilities. Our framework achieves state-of-the-art results on the AerialVLN-S dataset, demonstrating its effectiveness. In the future, we will explore flexible framework that will interact with humans to eliminate ambiguities in the instructions.

### v2

@@ A Simplified Input in the Cyclic Process
1. Lift off 5 meters. (Completed)
2. Turn 180 degrees. (Completed)
3. Head straight across the water to the road. (Completed)
4. Turn left to align the road. (In Process)
@@
IV Experiment
TABLE I: Comparison on the validation seen of the AerialVLN-S. * denotes we train these methods by running their source codes.
Validation Seen
model NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$
Random 109.6 0.0 0.0 0.0
Action Sampling 213.8 0.9 5.7 0.3
LingUNet [misra2018mapping] 383.8 0.6 6.9 0.2
Seq2Seq [anderson2018vision] 146.0 4.8 19.8 1.6
CMA [CMA] 121.0 3.0 23.2 0.6
AerialVLN [liu_2023_AerialVLN] 90.2 7.2 15.7 2.4
Navid* [navid] 105.1 6.8 15.5 1.1
Openfly* [openfly] 127.2 8.1 21.8 1.6
Ours 96.3 12.6 31.6 2.7
TABLE II: Comparison on the validation unseen of the AerialVLN-S. * denotes we train these methods by running their source codes.
Validation Unseen
model NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$
Random 149.7 0.0 0.0 0.0
Action Sampling 237.6 0.2 1.1 0.1
LingUNet [misra2018mapping] 368.4 0.4 3.6 0.9
Seq2Seq [anderson2018vision] 218.9 2.3 11.7 0.7
CMA [CMA] 172.1 3.2 16.0 1.1
AerialVLN [liu_2023_AerialVLN] 127.9 5.1 10.5 1.4
Navid* [navid] 106.9 6.1 12.3 0.7
Openfly* [openfly] 113.8 7.6 18.2 1.5
Ours 119.5 10.8 23.0 1.9
TABLE III: Comparison on OpenFly dataset unseen split.
OpenFly Unseen
model NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$
Random 301m 0.1 0.1 0
Seq2Seq[anderson2018vision] 291m 0.1 3.2 0.1
CMA[CMA] 268m 0.7 6.6 0.1
AerialVLN[liu_2023_AerialVLN] 227m 2.3 19.8 1.3
Navid[navid] 200m 2.6 16.9 2.2
OpenFly[openfly] 198m 5.1 27.3 2.0
Ours 205m 8.0 31.0 2.4
IV-A Dataset and Implementation Details
IV-A1 Dataset
We conduct experiments using two datasets to assess aerial VLN task performance, i.e., AerialVLN-S benchmark from [liu_2023_AerialVLN] and the recently introduced OpenFly dataset [openfly]. The AerialVLN-S dataset simulates real-world urban environments with over 870 object categories across diverse scenarios such as downtown cities, factories, parks, and villages. It contains 8,446 expert-curated flying paths recorded by AOPA-certified UAV pilots, ensuring trajectory realism and navigational fidelity. For evaluation, we utilize 12 scenes from its validation set and 5 scenes from the validation unseen set, as the official test set remains inaccessible. To enable broader comparative analysis, we introduce the OpenFly unseen dataset, which provides 1,200 test trajectories uniformly sampled from more diverse scenarios. We compare with several state-of-the-art VLN models on both benchmarks, using OpenFly’s test unseen trajectories for generalization assessment. This dual evaluation strategy rigorously tests the performances of these works, establishing comprehensive baselines for aerial VLN research.
IV-A2 Evaluation metrics
@@
To validate the effectiveness of our method, we utilize a comprehensive set of evaluation metrics following [liu_2023_AerialVLN]. We focus on several key aspects, i.e., Navigation Error (NE) quantifying the distance between the UAV’s stopping point and the actual destination, Success Rate (SR) which measures the proportion of navigations that successfully reach the destination within a 20-meter threshold, Oracle Success Rate (OSR), an idealized measure considering any point on the trajectory that comes within 20 meters of the destination as a success, and Success rate weighted by Normalised Dynamic Time Warping (SDTW), which takes into account both the navigation success rate and the similarity between the real paths.
IV-A3 Implementation details
@@
Our framework is implemented in both a simulator and a real environment. In the simulation environments of AirSim and Unreal Engine 4 (UE4), we evaluated the methods using a PC equipped with an Intel i9 12th-generation CPU and Nvidia RTX 4090 GPUs. For the real world environment, we test on a Q250 airframe, carrying an Intel RealSense D435i depth camera, and an NVIDIA Jetson Xavier NX running Ubuntu 18.04 as onboard computer. All methods are run on an external PC (same as the simulation setup) communicating with the onboard computer via a UDP network. For LLM reasoning we use the online API of GPT-4V and GPT-4o, where the default parameters are employed.
IV-B Experimental Results
IV-B1 Quantitative results in simulator
@@
Table I and  Table II compare our method with several other works on the AerialVLN-S dataset. Random and Action Sampling respectively mean sampling actions randomly and according to the action distribution of the training set. LingUNet [misra2018mapping] assumes an agent can see the destination from the beginning, which is transformed into a step-wise paradigm by [liu_2023_AerialVLN]. Seq2Seq [anderson2018vision] is a learning-based recurrent policy, directly predicting the next action. CMA is a baseline model in [liu_2023_AerialVLN], considering attention between multi-modalities. LAG is the main method in [liu_2023_AerialVLN], designing a look-ahead guidance to adjust paths. Navid [navid] is a VLM-based method, achieving great success in the indoor VLN task. OpenFly [openfly] proposes a keyframe-aware aerial VLN model, taking key historical observations into account. In comparison, our method significantly outperforms others in most metrics, demonstrating the superiority of our method. Notably, our method is training-free, while others are trained using corresponding data.  Table III shows the generalization results of different methods on the OpenFly Unseen dataset, where our method also achieves better performance than other works.
We found that instead of processing the entire long instruction and visual observations at each time step, the formatted decomposed sub-goal is more conducive to the fine-grained alignment of each action and landmark. Moreover, UAV has a better understanding of its own spatial position and other semantic constraints under the guidance of our STMR, which has obvious advantages in long instruction environments and leads to more robust execution. Aerial VLN is a challenging task that has not been fully studied, and the overall performance needs to be improved. We believe that our work can bring some insights into this field.
IV-B2 Quantitative results in real environment
@@
To test whether our solution works well in the real world, we collect 10 outdoor scenes including street scenes and forests with ground-truth lengths ranging from 50m to 500m. Then, we evaluate our method using a real UAV and call cloud-hosted LLMs to navigate in these challenging environments. Not only pretrained VLN models but also other LLM-based VLN methods, i.e., MapGPT [chen2024mapgpt] and NavGPT [zhou2024navgpt], are employed and compared in this experiment. Notably, MapGPT and NavGPT rely on predefined topological maps to predict the next navigation waypoint, we manually define candidate waypoints for them in this experiment. From demonstration snapshot Fig. 6, we can see that our method can align visual and textual landmarks as well as understand commands. Finally, the UAV reaches the destination successfully.  Table IV also shows that the proposed method achieves a much better performance than both pretrained VLN methods and training-free LLM-based methods. However, we can see that LLM-based methods including ours need more computation time. This is mainly because of the latency of the LLM API. We believe that with the development of LLMs and hardware computing power, the latency will gradually decrease.
Figure 4: Demonstration of the STMR in spatial reasoning.
Figure 5: Failure cases of our method.
TABLE IV: Quantitative results in real environment. The last column shows the average runtime of aerial VLN procedures.
model SR/$\uparrow$ OSR/$\uparrow$ Time (s)
AerialVLN [liu_2023_AerialVLN] 0 10 82
Navid [navid] 10 10 75
Openfly [openfly] 20 30 63
MapGPT [chen2024mapgpt] 20 20 176
NavGPT [zhou2024navgpt] 10 20 150
Ours 40 70 117
@@
Figure 6: Qualitative results in real environment. Short-range movements of the same action are merged into a single long-range movement, as shown in the image sequence.
IV-B3 Case analysis
Fig. 4 demonstrates the effectiveness of our STMR in helping spatial reasoning, where the left part shows an intermediate STMR during the navigation process, and the right part shows simplified prompt, questions, and reasoning answers. From Fig. 4 we can see that an LLM can understand and describe the surrounding environment through the STMR. It can even make accurate judgments about objects or areas at a distance. Such an ability is extremely helpful for aerial VLN in vast outdoor environments.
@@ IV-B3 Case analysis
We also summarize two most common failures caused by incorrect planning or execution. The first typical failure is the misunderstanding of ambiguous instructions. For example, as shown in the left part of Fig. 5, sometimes there are a lot of continuous commands without any landmark, such as “turn left, then move right, then go straight”, which have no objects for an LLM to refer to. This case will directly lead to the repeated execution of an action. The second typical error is caused by the inaccuracy of the visual perception. Although modern perception models show strong capability, they still need to be improved in identifying objects from different views. As a result, the key landmark may not be mapped to the STMR, causing the proper action not to be completed as illustrated in the right part of Fig. 5. This observation aligns with our real-world experiment.
TABLE V: Ablation study on different spatial representations.
Method Validation Unseen
NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$
Topo 203.3 4.9 12.8 1.8
Metric 165.0 6.1 13.7 2.2
Ours 87.1 19.4 26.2 5.7
IV-C Ablation Studies
@@
We carry out comprehensive ablation studies to assess the core designs of the proposed method. All ablation experiments are conducted on the randomly selected 100 samples from the AerialVLN-S dataset. Corresponding results are depicted in Table V,  Table VI,  Table VII, and Table VIII.
IV-C1 Different types of spatial information
@@
To demonstrate the spatial representation ability of STMR, we further compare it with other prompting formats as an LLM’s spatial information prompt. In Topo format (row 1 of Table V), we maintain a linguistically formed map that captures the topological relationships between different nodes. Each node records textual descriptions of visual observations, and the connectivity between nodes is described using textual prompts. For example, “Place 1 is connected with Places 2, 4, 0, 3”. In Metric format (row 2 of Table V), we arrange the visual observations from 8 different directions in a clockwise order relative to the agent’s current orientation and concatenate them into a single prompt. Each vision observation records the direction and distance of the landmark, for example, “a white building in the left front 10 meters away”. It is shown that row 2 improves SR by 1.2% (absolute) over row 1 but still has poorer navigation performance, since for some landmarks such as roads and rivers, UAVs still cannot obtain region-level spatial information based on a certain point distance description. Our STMR combining the semantic, topological, and metric information significantly enhances the UAV’s exploration capability and improves the accuracy in SR and OSR by over 10%.
IV-C2 Direct visual input vs STMR
@@
Due to the powerful multimodal representation capabilities of GPT-4V and GPT-4o, they are capable of direct reasoning with visual and textual prompts. However, it is challenging for them to make correct VLN decisions from sequential images. This may be because of the inability of VLMs in spatial-aware reasoning. As shown in Table VI, the proposed STMR representation with LLMs significantly outperforms the direct input of RGB images into VLMs.
IV-C3 Grid size of STMR
[…]
IV-C4 State update of sub-goal
To ensure that each sub-goal can be executed despite environmental changes, in each iteration, the LLM does not change the original path plan, but updates the status of the current sub-goal including todo, in process, and completed. We compare this strategy with updating the entire multi-step path planning (row 1 of Table VIII) according to the observation in each iteration. This strategy does not notably enhance the Oracle Success Rate (OSR), showing that the proposed STMR already furnishes enough capacity for navigating the environment comprehensively. Nonetheless, it positively impacts the decision-making process, resulting in an increase in the success rate (SR) from 9.0% to 10.8%.
TABLE VI: Ablation study on the different visual prompts for VLMs.
VLM Input RGB STMR NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$
GPT-4V $\checkmark$ 350.0 1.5 9.7 0.4
$\checkmark$ 193.1 5.6 17.4 0.9
GPT-4o $\checkmark$ 412.0 1.1 10.0 0.2
$\checkmark$ 119.5 10.8 23.0 2.6
TABLE VII: Ablation study on the grid size $r$ of STMR.
Size NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$
$10m$ 304.9 2.0 9.5 0.6
$5m$ 119.5 10.8 23.0 1.9
TABLE VIII: Ablation study on the state update strategies.
Module NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$
w/o state 289.7 9.0 22.1 1.7
w state 119.5 10.8 23.0 1.9
@@ V CONCLUSIONS AND FUTURE WORK
This paper addresses the challenging aerial VLN task by introducing an LLM-based training-free framework. To enhance the spatial reasoning ability of LLMs, we propose Semantic-Topo-Metric Representation (STMR). STMR first integrates instruction-related landmarks and their locations into a top-down map, and subsequently transforms this map into a matrix representation containing semantic, topological, and distance metric information. Taking the proposed STMR as a part of the LLM prompts, we significantly improve the UAV’s navigation capabilities. Our framework achieves state-of-the-art results on the AerialVLN-S and Openfly dataset, demonstrating its effectiveness. In the future, we will explore flexible framework that will interact with humans to eliminate ambiguities in the instructions.